how many grams of fe2o3 are needed to produce 51.9 g of cao


1
Expert's answer
2021-03-05T03:32:10-0500

Moles of Fe2O3 = 51.9 g / 159.69 g = 0.325moles

1 moles of Fe2O3 gives 1 moles of CaO

0.325 moles of Fe2O3 gives 0.325 moles of CaO

Gram of CaO = 0.325×56.077 = 18.22 g